Unit 26 Homework Assignment  Physics141-Summer 2009

Unit 26 homework after session 1

• Read chapter 30 in the textbook Understanding Physics.

• Work textbook problems 30-27, 30-29, 30-37. Unit 26 homework after session 2

• Read chapter 31 in the textbook Understanding Physics.

• Work textbook problems 31-5, 31-20, 31-29. Unit 26 homework after session 3

• Read chapter 32 in the textbook Understanding Physics.(sections 32-6, 32-7, 32-8, and 32-9 are optional)

• Work textbook problems 32-8, 32-22, 32-35.

All Unit 26 homework problems and the unit 26 activity guide are due on the first day of Unit 27